# Scalable magnetic resonance fingerprinting: Incremental inference of high dimensional elliptical mixtures from large data volumes

### v1

@@ Abstract
Magnetic Resonance Fingerprinting (MRF) is an emerging technology with the potential to revolutionize radiology and medical diagnostics. In comparison to traditional magnetic resonance imaging (MRI), MRF enables the rapid, simultaneous, non-invasive acquisition and reconstruction of multiple tissue parameters, paving the way for novel diagnostic techniques. In the original matching approach, reconstruction is based on the search for the best matches between in vivo acquired signals and a dictionary of high-dimensional simulated signals (fingerprints) with known tissue properties. A critical and limiting challenge is that the size of the simulated dictionary increases exponentially with the number of parameters, leading to an extremely costly subsequent matching. In this work, we propose to address this scalability issue by considering probabilistic mixtures of high-dimensional elliptical distributions, to learn more efficient dictionary representations. Mixture components are modelled as flexible ellipitic shapes in low dimensional subspaces. They are exploited to cluster similar signals and reduce their dimension locally cluster-wise to limit information loss. To estimate such a mixture model, we provide a new incremental algorithm capable of handling large numbers of signals, allowing us to go far beyond the hardware limitations encountered by standard implementations. We demonstrate, on simulated and real data, that our method effectively manages large volumes of MRF data with maintained accuracy. It offers a more efficient solution for accurate tissue characterization and significantly reduces the computational burden, making the clinical application of MRF more practical and accessible.
Dimension reduction,
Clustering,
Incremental learning,
High dimensional mixture models,
Elliptical distributions,
@@ 1 Introduction
Traditional Magnetic Resonance (MR) imaging relies on an analytical resolution of dynamical equations using conventional tuning of the MR hardware through sequences of pulses, each characterized by different values of parameters such as the flip angle and repetition time. Standard quantitative MRI (qMRI) methods are based on a single sequence for a single parameter measurement at a time. This leads to high scan times for multi-parametric protocols as each parameter estimate involves one MR sequence. A recent approach named Magnetic Resonance Fingerprinting (MRF, Ma et al. (2013)) has been developed to overcome these limitations. The MRF protocol involves fast undersampled acquisitions with time-varying parameters defining the MRF sequence that produces temporal signal evolutions (named fingerprints) in each voxel. In the original proposal, a dictionary search approach is used to compare the in vivo fingerprints with millions of numerical simulations of MR signals for which the associated parameters are known. These millions of simulated signals compose the so-called dictionary. The values of the parameters corresponding to the closest simulated signals or matches are then assigned to the associated in vivo voxels, allowing the simultaneous reconstruction of multiple quantitative maps (images) from extremely undersampled raw images, using only one single sequence, thus saving considerable acquisition time (Poorman et al., 2020; McGivney et al., 2020). From this, standard relaxometry MRF allows reconstructing parameter maps for relaxation times T1 and T2, over the whole human brain (1 mm3 spatial resolution) in 3 min (Ye et al., 2017; Gu et al., 2018) compared to 30 min for a standard T1/T2 exam. Moreover, the flexibility of the numerical simulations enables correction of system imperfections as well as some patient motions by including them in the model and post-processing pipelines (Bipin Mehta et al., 2019). Thus, MRF could be a game changer for emergency patients who need to complete exams in a few minutes. The power of the MRF approach is not limited to the estimation of relaxation times, in theory, it allows the measurement of any parameter that influences nuclear magnetization (e.g., microvascular networks), and could be added to the simulation model (Wang et al., 2019a; Coudert et al., 2024). However, increasing the number of estimated parameters, even moderately, induces the design of more complex sequences and increased reconstruction times, from hours to days. This limits the clinical application of high-dimensional MRF and necessitates the development of innovative processing methods. Consequently, a significant focus is on improving MRF reconstruction methods, as reviewed by Tippareddy et al. (2021) and Monga et al. (2024). In this work, we propose to focus on reducing the reconstruction times of MRF when more than the main two parameters, $T_{1}$, $T_{2}$, are involved, including the addition of $\delta\!f$ the frequency offset, the sensitivity of the magnetic field $B_{1}$, cerebral blood volume (CBV), and microvascular geometry (e.g., vessel radius denoted as $R$).
@@
MRF reconstruction is first recast as an inverse problem that can be solved using different approaches as recalled in the next section. All approaches make use, at some stage, of a dictionary of simulated pairs (parameters, signal), which represent our knowledge of the link between tissue parameters and MR time series, through a so-called direct or forward model. The dictionary is then either used to learn an inversion operator, from signal to tissue parameters, or to search for the best fits between observed signals and simulated ones. The approaches scalability relies thus greatly on their ability to extract efficiently the information encoded in simulations. Efficiency has different aspects: for search-type methods, the dictionary should not be too big, while for learning-based methods, the dictionary should be informative enough. These potentially opposite requirements call for efficient representations of simulated data. We propose to explore a divide-and-conquer strategy by introducing the framework of High Dimensional Mixtures of Elliptical Distributions (HD-MED). Probabilistic mixtures of high-dimensional elliptical distributions allow us to learn more efficient dictionary representations. Mixture components are modeled as flexible elliptical shapes in low-dimensional subspaces. They are exploited to cluster similar signals and reduce their dimension locally, at the cluster level, to limit information loss. To estimate such a mixture model, we provide a new incremental algorithm capable of handling large numbers of signals, allowing us to go far beyond the hardware limitations encountered by standard implementations. We demonstrate, on simulated and real data, that our method effectively manages large volumes of MRF data with maintained accuracy. It offers a more efficient solution for accurate tissue characterization and significantly reduces the computational burden, making the clinical application of MRF more practical and accessible.
@@ 2 MRF reconstruction as an inverse problem
An inverse problem refers to a situation where one aims to determine the causes of a phenomenon from experimental observations of its effects. In MRF, the goal is to infer from an observed signal a set of tissue characteristics or parameter values that best account for the signal. Such a resolution generally starts with modeling the phenomenon under consideration, which is called the ”direct” or ”forward problem.” It is generally assumed that at least the numerical evaluation of the forward model is available because experts have designed equations that can be solved either analytically or numerically. The most common use of the forward model is via a simulator that allows the creation of a database ${\cal D}_{f}$, usually referred to as a dictionary, of $N$ signals ${\mathbf{y}_{1},\ldots,\mathbf{y}_{N}}$ with $\mathbf{y}_{i}\in{\cal Y}\subset\mathbb{R}^{M}$, generated (stored or computed on the fly) by running the theoretical (physical) model $f$ for many different tissue parameter values ${\mathbf{t}_{1},\ldots,\mathbf{t}_{N}}$ with $\mathbf{t}_{i}\in{\cal T}\subset\mathbb{R}^{L}$ and $\mathbf{y}_{i}=f(\mathbf{t}_{i})$. The generated tissue parameter values then only partially represent the full space ${\cal T}$ of possible values and correspond to a discrete grid in the full space. In this context, we can distinguish two types of methods, referred to below as optimization and learning approaches.
@@ 2.1 Optimization vs learning (or regression) approaches
$\hat{\mathbf{t}}$$\in$ $\arg\min_{\mathbf{t}\in{\cal T}}d(\mathbf{y}_{obs},f,\mathbf{t})\;.$ (1)
Typically, $d(\mathbf{y}_{obs},f,\mathbf{t})=d(\mathbf{y}_{obs},f(\mathbf{t}))=||\mathbf{y}_{obs}-f(\mathbf{t})||^{2}\;.$ Solutions are searched in the full ${\cal T}$ space but solutions could also be penalized as done in grid search methods. Indeed, in grid search, the previous full search is replaced by a simpler look-up or matching operation making use of the database ${\cal D}_{f}$, often created beforehand offline. The search space is significantly reduced from a continuous space ${\cal T}$ to a discrete and finite ${\cal D}_{f}$. The speed gain is significant in comparison to traditional optimization methods as retrieving a value from memory is often faster than undergoing an expensive usually iterative computation. Their disadvantage is the instability of solutions. Many questions remain on how to choose the merit function, how many $\mathbf{y}_{n}$ in the look up table have to be kept to estimate parameters, how to choose the look-up table, etc. When the number of parameters is small, grid search is suitable and can provide very good predictions. However, for even moderate numbers of parameters, the required number of elements in the dictionary renders grid search either intractable or inaccurate. The technique is not amortized, for each new $\mathbf{y}_{obs}$, we have to compute the matching score $d(\mathbf{y}_{obs},\mathbf{y}_{n})$ for all $\mathbf{y}_{n}$ in the dictionary. When the dimension of $\mathbf{t}$ increases, the dictionary (and $N$) has to be larger too for better accuracy. The computation of $N$ matching scores can become time consuming.
@@
Regression or learning methods are more efficient in that sense and usually have better amortization properties. In contrast to the previous ones, this category of methods has the advantage of adapting easily to provide tractable solutions in the case of massive inversions of high-dimensional data. The main principle is to transfer the computational cost and time from individual pointwise predictions to the learning of a global inverse operator from ${\cal D}_{f}$. The advantage is that once the operator is learned, it can be used, at negligible cost, for very large numbers of new signals. Then, the nature of the inverse operator needs to be specified. Traditional learning or regression methods are not specifically designed for high dimensional data but there has been a large literature covering this case, see e.g. Giraud (2014) for a review. In MRF, a popular approach from Cohen, Zhu and Rosen (2018), uses a four-layer fully connected network (DRONE) to learn the dictionary signals for reconstructing $T_{1}$ and $T_{2}$ parameters. Even though DRONE provides good results, this perceptron-based model loses the temporal coherence of the signal. Recently, Cabini et al. (2024) proposed a Recurrent Neural Network (RNN) with long-short term memory (LSTM) blocks, which yields better reconstruction results for $T_{1}$ and $T_{2}$ with more robustness towards noisy acquisitions. However, when more parameters need to be estimated, such as vessel oxygenation or radius (Christen, Bolar and Zaharchuk, 2013; Christen et al., 2014), the dimensionality of the signals dramatically increases. Barrier et al. (2024) demonstrated that a simple RNN is prone to catastrophic forgetting, where the RNN well estimates the beginning of the signals but learns the end less effectively. To mitigate this issue, they proposed using bidirectional LSTM (bi-LSTM) blocks within the RNN architecture, ensuring that both the beginning and the end of the signal are efficiently learned by the new bi-LSTM blocks. Despite the good results in reconstructing $T_{1}$, $T_{2}$, and vascular parameters, bi-LSTM still faces challenges in learning the middle part of the signals.
@@
Generally, the dictionary is seen as a collection of simulated signals with no particular spatial correlation. Another way to simulate signals is to acquire real parametric maps from different subjects and then simulate the MRF images. This approach was first proposed by Soyak et al. (2021) and later improved by Gu et al. (2024). Both studies utilize a UNet (Ronneberger, Fischer and Brox, 2015) to infer $T_{1}$ and $T_{2}$ directly using the entire MRF image as an input to the network, preserving valuable spatial information. Given the high dimensionality of the signals, the authors proposed adding attention layers (Vaswani et al., 2017) to focus on the most important dimensions of the signals. More recently, Li and Hu (2024) highlighted the limitations of using CNN, which have a restricted receptive field and capture spatial information only locally. To overcome these limitations, the authors proposed using a Local-Global vision Transformer to capture spatial information globally as well. However, capturing spatial information has a cost. Indeed, one needs to acquire from a large group of subjects multiple $T_{1}$ and $T_{2}$ maps, which takes about 30 min for a complete exam, making the data acquisition process much more costly than using more conventional dictionaries. Additionally, since $T_{1}$ and $T_{2}$ need to be acquired at two different times, this method introduces more errors from the registration of the acquired maps.
@@
Finally, Boux et al. (2021) also proposed to use GLLiM (Deleforge, Forbes and Horaud, 2015), a model that casts MRF reconstruction into a Bayesian inverse problem and then solves it using a learning approach. This method takes into account the high-dimensional property of MRF signals by defining the low-dimensional variables as the regressors, which in our case are the tissue parameters. By doing so, they start learning the low-to-high regression model from which they can derive the forward model parameters and then the high-to-low regression model, from MRF signals to tissue parameters, as desired.
@@ 3 Related work and positioning
In practice, acquired MRF acquisitions come as 4D matrix, made of a time series of 3D MRI images where each voxel contains the acquired, potentially long, fingerprint signal. In this work, we focus on designing efficient representations of large highly precise grids of simulated signals counterparts.
3.1 Parcimonious representations of dictionaries
@@
The curse of dimensionality goes with what is often called the bless of dimensionality, which refers to the fact that in high dimensional data sets, useful information actually leaves in much smaller dimensional parts of the data space. One approach to the MRF reconstruction problem is then to reduce the dimension of the dictionary beforehand to reduce the matching or learning cost. The first to propose such an approach for efficient matching were McGivney et al. (2014), who applied Singular Value Decomposition (SVD) to the dictionary of signals. Once the decomposition is learned, one can project any new acquisition into the SVD low-dimensional subspace. However, when the size of the dictionary increases, computing the SVD becomes costly as it requires loading the complete dictionary into fast-access memory (e.g the RAM). To address this, Yang et al. (2018) proposed using randomized SVD (Halko, Martinsson and Tropp, 2011). Golbabaee et al. (2019) also suggested applying SVD to the dictionary before training a neural network. Other methods proposed a non-Euclidian analysis of the dictionary space projecting the signals into a lower-dimensional manifold (Li and Hu, 2023). Reducing the dimension of high-dimensional dictionaries assumes that most of the information in the signals can be captured and represented in a much lower-dimensional subspace. Classical techniques include principal component analysis (PCA, Jolliffe and Cadima (2016)), probabilistic principal component analysis (PPCA, Tipping and Bishop (1999a)), factor analyzers (FA), sparse models (Zou, Hastie and Tibshirani, 2006; d’Aspremont et al., 2007; Archambeau and Bach, 2008), and newer methods such as diffusion maps (Coifman and Lafon, 2006). More flexible approaches are based on mixtures of the previous ones, such as mixtures of factor analyzers (MFA, McLachlan, Peel and Bean (2003)) introduced by Ghahramani and Hinton (1997) and extended by McLachlan, Peel and Bean (2003); Baek and McLachlan (2011), and mixtures of PPCA (MPPCA, Tipping and Bishop (1999b); Xu, Balzano and Fessler (2023)) with recent generalizations (Hong et al., 2023; Xu, Balzano and Fessler, 2023). Another mixture approach is called HDDC in Bouveyron, Girard and Schmid (2007) and HD-GMM in Bouveyron and Brunet-Saumard (2014) for High Dimensional Gaussian Mixture Models, which encompass many forms of MFA and MPPCA and generalize them. In particular HD-GMM can be used to obtain multiple low-dimensional subspaces of different dimensions. For a review on high-dimensional clustering via mixtures, see Bouveyron and Brunet-Saumard (2014). Such a divide-and-conquer strategy has been used in MRF by first creating a global clustering of the signals and then reducing the dimension. This method was applied by Cauley et al. (2015), who performed K-way partitioning of the dictionary before using multiple cluster-associated PCA to reduce the dimension in each part. When matching a new signal, one first determines which cluster it belongs to and then applies the projection learned by the associated PCA model. More recently, Ullah et al. (2023) proposed a simpler approach: applying a clustering algorithm to the dictionary before utilizing GPUs for the dictionary matching enabling fast matching without the need for dimensionality reduction. However this method still faces issues when the number of parameters to estimate is larger than $3$.
@@
Most of these methods are designed for batch data and are thus sensitive to hardware limits such as memory, restricting the amount of data they can process. For instance, some dictionaries exceed terabyte sizes. A simple solution is to down-sample data sets before processing, potentially losing useful information. Another approach is to design incremental, also referred to as online, variants that handle data sequentially in smaller groups. A number of incremental approaches exist for dimension reduction techniques, see the recent SHASTA-PCA (Gilman et al., 2023) and references therein, or Balzano, Chi and Lu (2018) for a review. To our knowledge, much fewer solutions exist for mixtures. Estimation of such models is generally based on maximum likelihood estimation via the Expectation-Maximization (EM) algorithm (McLachlan and Krishnan (2008)). A preliminary attempt for an incremental MPPCA can be found in Bellas et al. (2013) but it is based on heuristic approximations of the EM steps.

### v2

# Scalable magnetic resonance fingerprinting: Incremental inference of high-dimensional elliptical mixtures from large data volumes

@@ Abstract
Magnetic Resonance Fingerprinting (MRF) is an emerging technology with the potential to revolutionize radiology and medical diagnostics. In comparison to traditional magnetic resonance imaging (MRI), MRF enables the rapid, simultaneous, non-invasive acquisition and reconstruction of multiple tissue parameters, paving the way for novel diagnostic techniques. In the original matching approach, reconstruction is based on the search for the best matches between in vivo acquired signals and a dictionary of high-dimensional simulated signals (fingerprints) with known tissue properties. A critical and limiting challenge is that the size of the simulated dictionary increases exponentially with the number of parameters, leading to an extremely costly matching. In this work, we propose to address this scalability issue by considering probabilistic mixtures of high-dimensional elliptical distributions, to learn more efficient dictionary representations. Mixture components are modelled as flexible elliptical shapes in low dimensional subspaces. They are exploited to cluster similar signals and reduce their dimension locally cluster-wise limiting information loss. To estimate such a mixture model, we provide a new incremental algorithm capable of handling large numbers of signals, allowing us to go far beyond the hardware limitations encountered by standard implementations. We demonstrate, on simulated and real data, that our method effectively manages large volumes of MRF data with maintained accuracy. It offers a more efficient solution for accurate tissue characterization and significantly reduces the computational burden, making the clinical application of MRF more practical and accessible.
@@ 1 Introduction
Traditional Magnetic Resonance (MR) imaging relies on an analytical resolution of dynamical equations using conventional tuning of the MR hardware through sequences of pulses, each characterized by different values of parameters such as the flip angle and repetition time. Standard quantitative MRI (qMRI) methods are based on a single sequence for a single parameter measurement at a time. This leads to high scan times for multi-parametric protocols as each parameter estimate involves one MR sequence. A recent approach named Magnetic Resonance Fingerprinting (MRF, ma2013) has been developed to overcome these limitations. The MRF protocol involves fast undersampled acquisitions with time-varying parameters defining the MRF sequence that produces temporal signal evolutions (named fingerprints) in each voxel. In the original proposal, a dictionary search approach is used to compare the in vivo fingerprints with millions of numerical simulations of MR signals for which the associated parameters are known. These millions of simulated signals compose the so-called dictionary. The values of the parameters corresponding to the closest simulated signals or matches are then assigned to the associated in vivo voxels, allowing the simultaneous reconstruction of multiple quantitative maps (images) from extremely undersampled raw images, using only one single sequence, thus saving considerable acquisition time (poorman2020; mcgivney2020). From this, standard relaxometry MRF allows reconstructing parameter maps for relaxation times T1 and T2, over the whole human brain (1 mm3 spatial resolution) in 3 min (ye2017; gu2018) compared to 30 min for a standard T1/T2 exam. Moreover, the flexibility of the numerical simulations enables correction of system imperfections as well as some patient motions by including them in the model and post-processing pipelines (bipin2019). Thus, MRF could be a game changer for emergency patients who need to complete exams in a few minutes. The power of the MRF approach is not limited to the estimation of relaxation times, in theory, it allows the measurement of any parameter that influences nuclear magnetization (e.g., microvascular networks), and could be added to the simulation model (Wang2019; coudert2024; coudert2025mr). However, increasing the number of estimated parameters, even moderately, induces the design of more complex sequences and increased reconstruction times, from hours to days. This limits the clinical application of high-dimensional MRF and necessitates the development of innovative processing methods. Consequently, a significant focus is on improving MRF reconstruction methods, as reviewed by tippareddy2021 and monga2024. In this work, we propose to focus on reducing the reconstruction times of MRF when more than the main two parameters, $T_{1}$, $T_{2}$, are involved, including the addition of $\delta\!f$ the frequency offset, the sensitivity of the magnetic field $B_{1}$, cerebral blood volume (CBV), and microvascular geometry (e.g., vessel radius denoted as $R$).
@@
MRF reconstruction is first recast as an inverse problem that can be solved using different approaches as recalled in the next section. All approaches make use, at some stage, of a dictionary of simulated pairs (parameters, signal), which represent our knowledge of the link between tissue parameters and MR time series, through a so-called direct or forward model. The dictionary is then either used to learn an inversion operator, from signal to tissue parameters, or to search for the best fits between observed signals and simulated ones. The approaches’ scalability relies thus greatly on their ability to extract efficiently the information encoded in simulations. Efficiency has different aspects: for search-type methods, the dictionary should not be too big, while for learning-based methods, the dictionary should be informative enough. These potentially opposite requirements call for efficient representations of simulated data. We propose to explore a divide-and-conquer strategy by introducing the framework of High Dimensional Mixtures of Elliptical Distributions (HD-MED). Probabilistic mixtures of high-dimensional elliptical distributions allow us to learn more efficient dictionary representations. Mixture components are modeled as flexible elliptical shapes in low-dimensional subspaces. They are exploited to cluster similar signals and reduce their dimension locally, at the cluster level, limiting information loss. To estimate such a mixture model, we provide a new incremental algorithm capable of handling large numbers of signals, allowing us to go far beyond the hardware limitations encountered with standard implementations. We demonstrate, on simulated and real data, that our method effectively manages large volumes of MRF data with maintained accuracy. It offers a more efficient solution for accurate tissue characterization and significantly reduces computational burden, making the clinical application of MRF more practical and accessible.
@@ 2 MRF reconstruction as an inverse problem
An inverse problem refers to a situation where one aims to determine the causes of a phenomenon based on experimental observations. In MRF, the goal is to infer a set of tissue characteristics or parameter values from an observed signal, identifying those that best explain the signal. Compared to standard MRI, which relies on sequential measurements with fixed parameters, MRF acquires data using a sequence of pseudo-random parameters (e.g., varying repetition times or flip angles). This process generates tissue-dependent time-series signals, where each time-series $\mathbf{y}$ - known as a fingerprint - is assigned to a specific 3D location (voxel) in the brain. Such a resolution generally starts with modeling the phenomenon under consideration, which is called the ”direct” or ”forward problem.” It is generally assumed that at least the numerical evaluation of the forward model is available because experts have designed equations that can be solved either analytically or numerically. The most common use of the forward model is via a simulator that allows the creation of a database ${\cal D}_{f}$, usually referred to as a dictionary, of $N$ signals $\mathbf{y}_{1},\ldots,\mathbf{y}_{N}$} with $\mathbf{y}_{i}\in{\cal Y}\subset\mathbb{R}^{M}$, generated (stored or computed on the fly) by running the theoretical (physical) model $f$ for many different tissue parameter values $\mathbf{t}_{1},\ldots,\mathbf{t}_{N}$} with $\mathbf{t}_{i}\in{\cal T}\subset\mathbb{R}^{L}$ and $\mathbf{y}_{i}=f(\mathbf{t}_{i})$. The typical dimension of each $\mathbf{y}_{i}$ considered in this paper is $M=260$. The generated tissue parameter values then only partially represent the full space $\cal T$} of possible values and correspond to a discrete grid in the space. In this context, we can distinguish two types of methods, referred to below as optimization and learning approaches.
@@ 2.1 Optimization vs learning (or regression) approaches
$\hat{\mathbf{t}} \displaystyle\in$ $\arg\min_{\mathbf{t}\in{\cal T}}d(\mathbf{y}_{obs},f,\mathbf{t})\;.$ (1)
[…]
@@
Regression or learning methods are more efficient in that sense and usually have better amortization properties. These methods have the advantage of providing tractable solutions in the case of massive inversions of high-dimensional data. The main principle is to transfer the computational cost and time from individual pointwise predictions to the learning of a global inverse operator from ${\cal D}_{f}$. The advantage is that once the operator is learned, it can be used, at negligible cost, for very large numbers of new signals. Then, the nature of the inverse operator needs to be specified. Traditional learning or regression methods are not specifically designed for high-dimensional data but there has been a large literature covering this case, see e.g. Giraud2014 for a review. In MRF, a popular approach from cohen2018, uses a four-layer fully connected network (DRONE) to learn the dictionary signals for reconstructing $T_{1}$ and $T_{2}$ parameters. Eventhough DRONE provides good results, this perceptron-based model loses the temporal coherence of the signal. Recently, cabini2024 proposed a Recurrent Neural Network (RNN) with long-short term memory (LSTM) blocks, which yields better reconstruction results for $T_{1}$ and $T_{2}$ with more robustness towards noisy acquisitions. However, when more parameters need to be estimated, such as vessel oxygenation or radius (christen2013; christen2014; coudert2025mr), the dimensionality of the signals dramatically increases. barrier2024 demonstrated that a simple RNN is prone to catastrophic forgetting, where the RNN well estimates the beginning of the signals but learns the end less effectively. To mitigate this issue, they proposed to use bidirectional LSTM (bi-LSTM) blocks within the RNN architecture, ensuring that both the beginning and the end of the signal are efficiently learned by the new bi-LSTM blocks. Despite the good results in reconstructing $T_{1}$, $T_{2}$, and vascular parameters, bi-LSTM still faces challenges in learning the middle part of the signals.
@@
Generally, the dictionary is seen as a collection of simulated signals with no particular spatial correlation. Another way to simulate signals is to acquire real parametric maps from different subjects and then simulate the MRF images. This approach was first proposed by soyak2021 and later improved by gu2024. Both studies utilize a UNet (ronneberger2015) to infer $T_{1}$ and $T_{2}$ directly using the entire MRF image as an input to the network, preserving valuable spatial information. Given the high-dimensionality of the signals, the authors proposed adding attention layers (vaswani2017) to focus on the most important dimensions of the signals. More recently, li2024 highlighted the limitations of using CNNs, which have a restricted receptive field and capture spatial information only locally. To overcome these limitations, the authors proposed using a Local-Global vision Transformer to capture spatial information globally as well. However, capturing spatial information has a cost. Indeed, one needs to acquire, from a large group of subjects, multiple $T_{1}$ and $T_{2}$ maps, which takes about 30 min for a complete exam, making the data acquisition process much more costly than using more conventional dictionaries. Additionally, since $T_{1}$ and $T_{2}$ need to be acquired at two different times, this method introduces more errors from the registration of the acquired maps.
@@
Finally, boux2021 also proposed to use GLLiM (Deleforge2015), a model that casts MRF reconstruction into a Bayesian inverse problem and then solves it using a learning approach. This method takes into account the high-dimensional property of MRF signals by defining the low-dimensional variables as the regressors, which in our case are the tissue parameters. By doing so, they start learning the low-to-high regression model from which they can derive the forward model parameters and then the high-to-low regression model, from MRF signals to tissue parameters, as desired.
@@ 3 Related work and positioning
In practice, acquired MRF acquisitions come as a 4D matrix, made of a time series of 3D MRI images where each voxel contains the acquired, potentially long, fingerprint signal. In this work, we focus on designing efficient representations of large highly precise grids of simulated signals counterparts.
3.1 Parsimonious representations of dictionaries
@@
The curse of dimensionality goes with what is often called the bless of dimensionality, which refers to the fact that in high-dimensional data sets, useful information actually lives in much smaller-dimensional parts of the data space. An approach to the MRF reconstruction problem is then to reduce the dimension of the dictionary beforehand to reduce the matching or learning cost. The first to propose such an approach for efficient matching were mcgivney2014, who applied Singular Value Decomposition (SVD) to the dictionary of signals. Once the decomposition is learned, one can project any new acquisition into the SVD low-dimensional subspace. However, when the size of the dictionary increases, computing the SVD becomes costly as it requires loading the complete dictionary into fast-access memory (e.g. the RAM). To address this, yang2018 proposed using randomized SVD (halko2011). golbabaee2019 also suggested applying SVD to the dictionary before training a neural network. Other methods proposed a non-Euclidean analysis of the dictionary space projecting the signals into a lower-dimensional manifold (li2023). Reducing the dimension of high-dimensional dictionaries assumes that most of the information in the signals can be captured and represented in a much lower-dimensional subspace. Classical techniques include principal component analysis (PCA, jolliffe2016), probabilistic principal component analysis (PPCA, tipping1999ppca), factor analyzers (FA) (Freitas2004), sparse models (Zou2006; daspremont2007; archambeau2008a), and newer methods such as diffusion maps (coifman2006). More flexible approaches are based on mixtures of the previous ones, such as mixtures of factor analyzers (MFA, mclachlan2003) introduced by gharamanietal97; gharamanietal99 and extended by mclachlan2003; baek2011mixtures, and mixtures of PPCA (MPPCA, tipping1999mppca; xu2023) with recent generalizations (hong2023; xu2023). Another mixture approach is called HDDC in bouveyron2007 and HD-GMM in bouveyron2014 for High Dimensional Gaussian Mixture Models, which encompass many forms of MFA and MPPCA and generalize them. In particular, HD-GMM can be used to obtain multiple low-dimensional subspaces of different dimensions. For a review on high-dimensional clustering via mixtures, see bouveyron2014. Such a divide-and-conquer strategy has been used in MRF by first creating a global clustering of the signals and then reducing the dimension. This method was applied by cauley2015, who performed K-way partitioning of the dictionary before using multiple cluster-associated PCA to reduce the dimension in each part. When matching a new signal, one first determines which cluster it belongs to and then applies the projection learned by the associated PCA model. More recently, ullah2023 proposed a simpler approach: applying a clustering algorithm to the dictionary, before utilizing GPUs for the dictionary matching, enabling fast matching without the need for dimensionality reduction. However, this method still faces issues when the number of parameters to estimate is larger than $3$.
@@
Most of these methods are designed for batch data and are thus sensitive to hardware limits such as memory, restricting the amount of data they can process. For instance, some dictionaries exceed terabyte (TB) sizes. A simple solution is to down-sample data sets before processing, potentially losing useful information. Another approach is to design incremental, also referred to as online, variants that handle data sequentially in smaller groups. Although such approaches can be used for genuine streaming data that is not stored after being seen, in this work we consider cases where past samples can possibly be revisited. A number of incremental approaches exist for dimension reduction techniques, see the recent SHASTA-PCA (gilman2023bis) and references therein, or balzano2018 for a review. To our knowledge, far fewer solutions exist for mixtures. Estimation of such models is generally based on maximum likelihood estimation via the Expectation-Maximization (EM) algorithm (mcLachlan08). A preliminary attempt for an incremental MPPCA can be found in bellas2013 but it is based on heuristic approximations of the EM steps.